Use substitution to determine if the given ordeded pair is an element of the solution set ofr the system of equations

(2,8)
y-2x=4
2y+x=-8

Found 2 solutions by rfer, richard1234:
Answer by rfer(16322)

y-2*2=4
y=8
-------------
2*8+x=-8
16+x=-8
x=24
doesn't work

Answer by richard1234(7193)

The ordered pair doesn't satisfy the second equation.